# Food as Soft Power: Taiwanese Gastrodiplomacy on Social Media and Algorithmic Suppression

### v1

@@ Introduction
On January 19, 2023, Google commemorated bubble tea with a caption stating, “Originating in Taiwan, this cold drink usually consists of tea, milk, and tapioca pearls. Bubble tea gained such popularity globally that it [is] officially announced as a new emoji on this day” (Google Doodles 2023). Since 2010, bubble tea has achieved global prominence, with Taiwanese immigrants establishing shops in Hong Kong, the United States, and the United Kingdom, and gradually in Dubai, Manila, and Mumbai—even in regions with smaller Taiwanese diaspora communities (Taiwan Today 2011). Given Taiwan’s contentious history with China, the increasing visibility of bubble tea has taken on symbolic significance, serving as a marker of Taiwanese sovereignty both physically and, increasingly, in digital spaces.
@@
Globally, news consumption increasingly occurs online, and according to Pew Research, more than half of young American adults—54%—now obtain their news from social media (Pew Research Center 2024). Instagram, YouTube, and Facebook are the most influential platforms, ranked first, second, and third, respectively, for social media news consumption. Visual-first platforms appear especially well-suited for promoting lifestyle and culinary content, facilitating affective connections between users and cultures. This raises an important question: what role can social media play in advancing soft power? Such potential is not without precedent—platforms have been weaponized to spread misinformation that influences elections and international relations (Chang and Fang 2024). In response, companies such as Meta (owner of Facebook and Instagram) have sought to reduce users’ exposure to political content in an effort to limit misinformation and its effects (Treisman 2024). However, what is deemed “political” can have direct consequences for Taiwan—an inherently political subject— thereby complicating the relationship between platform policies and cultural diplomacy.
@@
Gastrodiplomacy has long been a central component of Taiwan’s public diplomacy. Multiple branches of government have invested in promoting Taiwan’s culinary culture abroad, ranging from the Executive Yuan’s NT\$1.1 billion (approximately US\$34.08 million) investment between 2010 and 2013 to support the international expansion of the food service sector (Taiwan Today 2010) to President Lai Ching-te’s 2024 initiative to elevate Taiwan’s global profile by showcasing local culinary and tourism assets through cross-regional collaborations and targeted marketing campaigns (Radio Taiwan International 2024).
@@
By contrast, countries such as France and Japan have long leveraged culinary diversity as a soft power asset. Japan promotes a wide array of dishes—including ramen, sushi, and yakitori—through initiatives such as the global “Cool Japan/Global Washoku” campaign launched in 2006 (Cabral et al. 2024). French cuisine is similarly central to its national image, from the ubiquitous presence of French restaurants worldwide to the Elysée Palace’s use of haute cuisine in state dinners (Barroux 2022). French gastronomy has even been inscribed on UNESCO’s Intangible Cultural Heritage Lists since 2010 (Ministère de l’Europe et des Affaires Étrangères 2010). More recently, South Korea has pursued a synergistic strategy, combining its global entertainment exports (the “Hallyu” wave) with the promotion of traditional dishes such as kimchi, ramyun, tteokbokki, and bibimbap (Nihayati, Pertiwi, and Haripa 2022). Government initiatives have even enlisted K-pop groups to promote Korean cuisine internationally (Mishan 2022). Such examples highlight how culinary diversity can enhance gastrodiplomacy and raise questions about the relative importance of variety and cultural breadth in sustaining soft power.
@@
The rise of social media has extended the reach of Taiwanese gastrodiplomacy into digital arenas. While Taiwanese cuisine shares many techniques, ingredients, and flavor profiles with Chinese and Japanese cooking (Sui 2015), the Taiwanese government has consistently framed its culinary heritage as a unique cultural asset warranting promotion. Social media now serves as a transnational public sphere capable of diffusing this distinct food culture. Indeed, a large-scale study of over 1.8 billion English-language Facebook posts and 51 million Chinese-language Weibo posts found that online conversations reflect the core–periphery structure predicted by World Systems Theory: international relations are shaped not only by news coverage but also by patterns of public attention, which in turn influence perceptions of foreign affairs (Barnett et al. 2017). In this sense, social media platforms could serve as vital channels for promoting Taiwanese food culture and, by extension, advancing Taiwan’s soft power.
One notable instance of organic gastrodiplomacy is the emergence of the Milk Tea Alliance in 2020, originally involving Taiwan, Hong Kong, and Thailand—each with its own distinct milk tea tradition. The alliance originated from online disputes with Chinese users and bots, eventually coalescing into the hashtag #MilkTeaAlliance as a means of uniting voices opposing China’s political actions: contesting its claims over Taiwan, resisting Hong Kong’s crackdown under the National Security Law, and addressing political grievances in Thailand (Schaffar and Praphakorn 2021). For Taiwan, the movement coincided with the aftermath of the 2020 presidential election, when sovereignty and self-determination were central public concerns. Similar cross-regional solidarity had surfaced earlier in Hong Kong’s 2014 Sunflower Movement, encapsulated in the slogan “Today’s Hong Kong, Tomorrow Taiwan,” which gained renewed prominence after the 2020 security law (Chang, Haider, and Ferrara 2021). Even after Hong Kong’s democratic freedoms were curtailed, solidarity networks persisted, including “Milk Tea Alliance – Japan” and the Taiwan Digital Diplomacy Association’s “Milk Tea Alliance Project,” designed to promote democracy and awareness of Taiwan. These developments demonstrate how bubble tea, as a symbol, became intertwined with democratic values through international digital activism.
@@
Exposure refers to the visibility of content in users’ feeds, shaped by algorithms that evaluate factors such as engagement history, user preferences, content relevance, and recency (Gunina 2023). Engagement encompasses user interactions—most prominently, likes and comments—that signal content value. Likes are relatively low-effort actions that indicate approval and may enhance visibility indirectly, whereas comments represent a higher level of engagement that can more strongly influence algorithmic ranking (Rockower 2011). Algorithmic amplification occurs when platforms’ machine learning systems prioritize certain posts based on these signals, often increasing the reach of content with higher engagement. However, such processes can inadvertently favor sensational or misleading information, contributing to the formation of echo chambers that limit exposure to diverse perspectives (Flaxman, Goel, and Rao 2016).
@@
Historical cases illustrate both the democratizing potential and the risks of algorithm-driven communication. During the Arab Spring, social media facilitated mobilization by enabling citizens to bypass state-controlled media, coordinate mass gatherings, and broadcast events to global audiences (Wolfsfeld, Segev, and Sheafer 2013; Smidi and Shahin 2017). In Egypt, for example, once users circumvented internet restrictions, there was a surge in tweets tagged #egypt, connecting domestic and international conversations on political developments (Bruns, Highfield, and Burgess 2013). Similarly, the George Floyd protests in 2020 saw the hashtag #JusticeforGeorgeFloyd go viral, with widespread participation on Twitter and Facebook from diverse global audiences, further amplifying the movement’s visibility (Chang, Richardson, and Ferrara 2022). These cases highlight social media’s potential to advance democratic movements through rapid, decentralized information sharing.
@@
However, the same dynamics can exacerbate polarization and information manipulation. The 2016 U.S. presidential election and the Brexit referendum demonstrated how foreign actors—most notably from Russia and China— used bots, disinformation campaigns, and targeted narratives to deepen social divisions (Francia 2018; Del Vicario et al. 2017; Ferrara et al. 2020). Domestically, misinformation regarding COVID-19 mask usage and the integrity of mail-in ballots, largely circulated within conservative networks, influenced political discourse during the 2020 Democratic primaries (Chen et al. 2021).
@@
In response to such concerns, Meta collaborated with academic researchers to examine algorithmic influence during the 2020 U.S. presidential election. Findings suggest that altering exposure to different types of information had minimal impact on users’ political opinions or polarization levels, underscoring the challenges of influencing attitudes within entrenched echo chambers (Nyhan et al. 2023). Meta’s Third-Party Fact-Checking Program further revealed that misinformation was disproportionately concentrated in conservative groups and was especially prevalent on Facebook News (González-Bailón et al. 2023). Additional studies show that reducing social media usage ahead of elections did not significantly alter polarization or political views, though it reduced the likelihood of questioning misleading content (Allcott et al. 2024). Even presenting content in reverse-chronological order—thereby increasing exposure to political and low-credibility content— had little measurable effect on political knowledge or attitudes (Guess et al. 2023).
@@
Basic User Engagement via Likes of Bubble Tea Posts (English versus Chinese)
@@
To investigate the key topics in the dataset, we utilized social network analysis, time series analysis, and Latent Dirichlet Allocation (LDA) for a robustness check (Appendix D). We also counted the most common food types to identify prominent items, potential connections, and latent themes within the hashtags and text sections of the posts that might affect the algorithm and engagement metrics (Blei, Ng, and Jordan 2003; Jelodar et al. 2019). This approach allowed for visualizations and robust insight checks into the semantic space, where, at the word level, we identify key themes and their relationships. Crucially, this process helps reveal clusters and patterns in which social or political ideas may intersect with Taiwanese food.
@@
To systematically compare the representation of food types across similar countries and evaluate their potential use in gastrodiplomacy, we computed the Shannon entropy for food-type mentions in online discourse using relative frequencies extracted from Instagram datasets over the same period as those analyzed for Taiwan. Shannon entropy provides a quantitative measure of diversity: higher values indicate a more even, heterogeneous spread among categories, while lower values signify concentration around a few dominant items.
@@ Results
Table 1 shows the distribution of food hashtags and their counts, sorted by frequency. Bubble tea is overwhelmingly dominant, appearing more than 95% of the time. In contrast, soup dumplings occupy only 2% and beef noodle soup takes up 0.66%, which marks a precipitous drop from bubble tea and milk tea’s dominance. Notably, this distribution resembles a power-law distribution, which is common among social systems that exhibit the rich-get-richer phenomenon (Clauset, Shalizi, and Newman 2009).
@@
Figure 4 shows the time series of bubble tea mentions over the past decade on Instagram. There is a clear seasonality in the attention given to bubble tea. While this may partly be due to increased consumption of cold drinks during the summer, the inciting factor is the Milk Tea Alliance in early 2020, which persisted throughout the summer as the Hong Kong protests continued. This suggests that broader socio-political events can drive heightened online attention to bubble tea. Additionally, peaks emerge annually around National Bubble Tea Day (April 30), an observance established by the Taiwanese-American chain Kung Fu Tea in 2018 (Kung Fu Tea 2025). These recurring spikes indicate a correlation between targeted promotions or social movements and an increase in bubble tea-related posts.
@@
Figure 5 shows the time series of bubble tea metrics, providing further evidence of the polarizing nature of the first summer with bubble tea. Specifically, it displays the standardized time series of posts (blue), likes (orange), comments (green), and views (red). Compared to all subsequent summers, the first summer features a significant bifurcation in comments and likes. In contrast, the summer of 2022 has many more likes relative to comments. The ratio between these two metrics matters: in social media studies, “rationing” refers to the phenomenon where the ratio of comments to likes can signal the level of engagement or controversy surrounding a post (Minot et al. 2021; Riedl, Lukito, and Woolley 2023). A higher comment-to-like ratio often indicates a more polarizing or contentious topic. Therefore, the elevated comments seen in 2020 likely reflect the broader socio-political tensions at the time—especially the Hong Kong protests and Tsai’s 2020 inauguration—prompting users to voice their opinions more actively than simply liking the posts.

### v2

@@ Introduction
On January 19, 2023, Google commemorated bubble tea with a caption stating, “Originating in Taiwan, this cold drink usually consists of tea, milk, and tapioca pearls. Bubble tea gained such popularity globally that it [is] officially announced as a new emoji on this day” (Google2023). Since 2010, bubble tea has achieved global prominence, with Taiwanese immigrants establishing shops in Hong Kong, the United States, and the United Kingdom, and gradually in Dubai, Manila, and Mumbai—even in regions with smaller Taiwanese diaspora communities (TaiwanToday2011). Given Taiwan’s contentious history with China, the increasing visibility of bubble tea has taken on symbolic significance, serving as a marker of Taiwanese sovereignty both physically and, increasingly, in digital spaces.
@@
Globally, news consumption increasingly occurs online, and according to Pew Research, more than half of young American adults—54%—now obtain their news from social media (Pew2024). Instagram, YouTube, and Facebook are the most influential platforms, ranked first, second, and third, respectively, for social media news consumption. Visual-first platforms appear especially well-suited for promoting lifestyle and culinary content, facilitating affective connections between users and cultures. This raises an important question: what role can social media play in advancing soft power? Such potential is not without precedent—platforms have been weaponized to spread misinformation that influences elections and international relations (Chang2024; ChangWangFang2024). In response, companies such as Meta (owner of Facebook and Instagram) have sought to reduce users’ exposure to political content in an effort to limit misinformation and its effects (Treisman2024). However, what is deemed “political” can have direct consequences for Taiwan—an inherently political subject— thereby complicating the relationship between platform policies and cultural diplomacy.
@@
Gastrodiplomacy has long been a central component of Taiwan’s public diplomacy. Multiple branches of government have invested in promoting Taiwan’s culinary culture abroad, ranging from the Executive Yuan’s NT\$1.1 billion (approximately US\$34.08 million) investment between 2010 and 2013 to support the international expansion of the food service sector (TaiwanToday2010) to President Lai Ching-te’s 2024 initiative to elevate Taiwan’s global profile by showcasing local culinary and tourism assets through cross-regional collaborations and targeted marketing campaigns (RTI2024).
@@
By contrast, countries such as France and Japan have long leveraged culinary diversity as a soft power asset. Japan promotes a wide array of dishes—including ramen, sushi, and yakitori—through initiatives such as the global “Cool Japan/Global Washoku” campaign launched in 2006 (Cabral2024). French cuisine is similarly central to its national image, from the ubiquitous presence of French restaurants worldwide to the Elysée Palace’s use of haute cuisine in state dinners (Barroux2022). French gastronomy has even been inscribed on UNESCO’s Intangible Cultural Heritage Lists since 2010 (MEAE2010). More recently, South Korea has pursued a synergistic strategy, combining its global entertainment exports (the “Hallyu” wave) with the promotion of traditional dishes such as kimchi, ramyun, tteokbokki, and bibimbap (Nihayati2022). Government initiatives have even enlisted K-pop groups to promote Korean cuisine internationally (Mishan2022). Such examples highlight how culinary diversity can enhance gastrodiplomacy and raise questions about the relative importance of variety and cultural breadth in sustaining soft power.
@@
The rise of social media has extended the reach of Taiwanese gastrodiplomacy into digital arenas. While Taiwanese cuisine shares many techniques, ingredients, and flavor profiles with Chinese and Japanese cooking (Sui2015), the Taiwanese government has consistently framed its culinary heritage as a unique cultural asset warranting promotion. Social media now serves as a transnational public sphere capable of diffusing this distinct food culture. Indeed, a large-scale study of over 1.8 billion English-language Facebook posts and 51 million Chinese-language Weibo posts found that online conversations reflect the core–periphery structure predicted by World Systems Theory: international relations are shaped not only by news coverage but also by patterns of public attention, which in turn influence perceptions of foreign affairs (Barnett2017). In this sense, social media platforms serve as vital channels for promoting Taiwanese food culture and, by extension, advancing Taiwan’s soft power. Crucially, due to its shared history with China and alliance with the United States, Taiwan serves as an information buffer between autocratic and democratic states (weener2025beyond).
One notable instance of organic gastrodiplomacy is the emergence of the Milk Tea Alliance in 2020, originally involving Taiwan, Hong Kong, and Thailand—each with its own distinct milk tea tradition. The alliance originated from online disputes with Chinese users and bots, eventually coalescing into the hashtag #MilkTeaAlliance as a means of uniting voices opposing China’s political actions: contesting its claims over Taiwan, resisting Hong Kong’s crackdown under the National Security Law, and addressing political grievances in Thailand (Schaffar2021). For Taiwan, the movement coincided with the aftermath of the 2020 presidential election, when sovereignty and self-determination were central public concerns. Similar cross-regional solidarity had surfaced earlier in Hong Kong’s 2014 Sunflower Movement, encapsulated in the slogan “Today’s Hong Kong, Tomorrow Taiwan,” which gained renewed prominence after the 2020 security law (Chang2021). Even after Hong Kong’s democratic freedoms were curtailed, solidarity networks persisted, including “Milk Tea Alliance – Japan” and the Taiwan Digital Diplomacy Association’s “Milk Tea Alliance Project,” designed to promote democracy and awareness of Taiwan. These developments demonstrate how bubble tea, as a symbol, became intertwined with democratic values through international digital activism.
@@
Exposure refers to the visibility of content in users’ feeds, shaped by algorithms that evaluate factors such as engagement history, user preferences, content relevance, and recency (Gunina2023). Engagement encompasses user interactions—most prominently, likes and comments—that signal content value. Likes are relatively low-effort actions that indicate approval and may enhance visibility indirectly, whereas comments represent a higher level of engagement that can more strongly influence algorithmic ranking (Rockower2011). Algorithmic amplification occurs when platforms’ machine learning systems prioritize certain posts based on these signals, often increasing the reach of content with higher engagement. However, such processes can inadvertently favor sensational or misleading information, contributing to the formation of echo chambers that limit diverse perspectives (Flaxman2016).
@@
Historical cases illustrate both the democratizing potential and the risks of algorithm-driven communication. During the Arab Spring, social media facilitated mobilization by enabling citizens to bypass state-controlled media, coordinate mass gatherings, and broadcast events to global audiences (Wolfsfeld2013; Smidi2017). In Egypt, for example, once users circumvented internet restrictions, there was a surge in tweets tagged #egypt, connecting domestic and international conversations on political developments (Bruns2013). Similarly, the George Floyd protests in 2020 saw the hashtag #JusticeforGeorgeFloyd go viral, with widespread participation on Twitter and Facebook from diverse global audiences, further amplifying the movement’s visibility (Chang2022). These cases highlight social media’s potential to advance democratic movements through rapid, decentralized information sharing.
@@
However, the same dynamics can exacerbate polarization and information manipulation. The 2016 U.S. presidential election and the Brexit referendum demonstrated how foreign actors—most notably from Russia and China— used bots, disinformation campaigns, and targeted narratives to deepen social divisions (Francia2018; DelVicario2017; Ferrara2020). Domestically, misinformation regarding COVID-19 mask usage and the integrity of mail-in ballots, largely circulated within conservative networks, influenced political discourse during the 2020 Democratic primaries (Chen2021).
@@
In response to such concerns, Meta collaborated with academic researchers to examine algorithmic influence during the 2020 U.S. presidential election. Findings suggest that altering exposure to different types of information had minimal impact on users’ political opinions or polarization levels, underscoring the challenges of influencing attitudes within entrenched echo chambers (Nyhan2023). Meta’s Third-Party Fact-Checking Program further revealed that misinformation was disproportionately concentrated in conservative groups and was prevalent on Facebook News (Gonzalez2023). Additional studies show that reducing social media usage ahead of elections did not significantly alter polarization or political views, though it reduced the likelihood of questioning misleading content (Allcott2024). Even presenting content in reverse-chronological order—thereby increasing exposure to political and low-credibility content— had little measurable effect on political knowledge or attitudes (Guess2023).
@@
To investigate the key topics in the dataset, we utilized social network analysis, time series analysis, and Latent Dirichlet Allocation (LDA) for a robustness check (Appendix D). We also counted the most common food types to identify prominent items, potential connections, and latent themes within the hashtags and text sections of the posts that might affect the algorithm and engagement metrics (Blei2003LDA; Jelodar2019TopicModeling). This approach allowed for visualizations and robust insight checks into the semantic space, where, at the word level, we identify key themes and their relationships. Crucially, this process helps reveal clusters and patterns in which social or political ideas may intersect with Taiwanese food.
@@
To systematically compare the representation of food types across similar countries and evaluate their potential use in gastrodiplomacy, we computed the Shannon entropy for food-type mentions in online discourse using relative frequencies extracted from Instagram datasets over the same period as those analyzed for Taiwan. Shannon entropy provides a quantitative measure of diversity frequently used in social science (chang2025liberals). Higher values indicate a more even, heterogeneous spread among categories, while lower values signify concentration around a few dominant items.
@@ Results
Table 1 shows the distribution of food hashtags and their counts, sorted by frequency. Bubble tea is overwhelmingly dominant, appearing more than 95% of the time. In contrast, soup dumplings occupy only 2% and beef noodle soup takes up 0.66%, which marks a precipitous drop from bubble tea and milk tea’s dominance. Notably, this distribution resembles a power-law distribution, which is common among social systems that exhibit the rich-get-richer phenomenon (clauset2009powerlaw).
@@
Figure 4 shows the time series of bubble tea mentions over the past decade on Instagram. There is a clear seasonality in the attention given to bubble tea. While this may partly be due to increased consumption of cold drinks during the summer, the inciting factor is the Milk Tea Alliance in early 2020, which persisted throughout the summer as the Hong Kong protests continued. This suggests that broader socio-political events can drive heightened online attention to bubble tea. Additionally, peaks emerge annually around National Bubble Tea Day (April 30), an observance established by the Taiwanese-American chain Kung Fu Tea in 2018 (kungfu2025). These recurring spikes indicate a correlation between targeted promotions or social movements and an increase in bubble tea-related posts.
@@
Figure 5 shows the time series of bubble tea metrics, providing further evidence of the polarizing nature of the first summer with bubble tea. Specifically, it displays the standardized time series of posts (blue), likes (orange), comments (green), and views (red). Compared to all subsequent summers, the first summer features a significant bifurcation in comments and likes. In contrast, the summer of 2022 has many more likes relative to comments. The ratio between these two metrics matters: in social media studies, “rationing” refers to the phenomenon where the ratio of comments to likes can signal the level of engagement or controversy surrounding a post (minot2021; riedl2023). A higher comment-to-like ratio often indicates a more polarizing or contentious topic. Therefore, the elevated comments seen in 2020 likely reflect the broader socio-political tensions at the time—especially the Hong Kong protests and Tsai’s 2020 inauguration—prompting users to voice their opinions more actively than simply liking the posts.